# Intergenerational Equity of Climate Change Mitigation

### v1

@@ Methods
where $C_{\mathrm{A}}(t)$ and $C_{\mathrm{D}}(t)$ denote the abatement cost and damage cost, respectively.
Equation 2 allows us to interpret a change in a parameter of the abatement policy as a change in the cash-flow structure of a financial product. Here $\mathrm{d}t(C_{\mathrm{A}}(t)+C_{\mathrm{D}}(t))$ corresponds to the time-$t$ cash-flow, $\frac{\partial V}{\partial C(t)}$ to a weight (like a notional) and $\frac{N(0)}{N(t)}$ to the discount factor.
From eq. 2, we see that the optimisation across generations may result in intergenerational inequality: Since $\mathrm{d}C_{\mathrm{D}}/\mathrm{d}T^{\mu=1}>0$ and $\mathrm{d}C_{\mathrm{A}}/\mathrm{d}T^{\mu=1}<0$, the calibration is equating the marginal cost changes of abatement and damage. However, these changes occur at different times, maybe even at different generations. Note that eq. 2 implies eq. 1.
Details on different extensions of DICE-2016-R
We endowed the classical DICE model with several extensions. These are not specific to the DICE model and can also be applied to other IAMs. They are also independent and parameterised in a way that allows to recover the classical model for a specific parameter choice.
Stochastic Discount Rates
The interest rate $r$ is modelled with a classical interest rate model. The results depicted here were produced with a Hull-White model with a constant mean-reversion and volatility. The implementation allows to replace this model with a general term structure model.
Stochastic Abatement Policy
It is well understood that the optimal abatement policy depends on the interest rate level. Hence, postulating stochastic interest rates, the abatement policy should be allowed to adapt to the interest rate level. This can also be interpreted as illustrating the effect of a repeated re-calibration of the model to changed interest rate levels. We assume a simple, functional model that depends on the spot interest rate, namely
@@
The simplicity of the models allows us to interpret the parameters. We have that $a_{1}$ in (3) describes the change of the abatement speed by interest rate change,
$a_{1}\ =\ \frac{\partial}{\partial r}\ \frac{\partial\mu(t)}{\partial t}\text{.}$
@@
Obviously, the stochastic abatement models may achieve a larger-or-equal welfare or lower-or-equal welfare-risk than the corresponding deterministic model simply due to the additional degree of freedom $a_{1}$ in the abatement optimisation.
Funding of Abatement Cost
Let $C_{\mu}(t)$ denote the cost that occur in time $t$ due to the abatement policy $\mu$. In the classical model, the total cost that occur in time $t$ are
$C(t)\ :=\ C_{\mathrm{A}}(t)+C_{\mathrm{D}}(t)\text{.}$
where $C_{\mathrm{D}}(t)$ are the damage costs and $C_{\mathrm{A}}(t)=C_{\mu}(t)$ are the abatement costs.
We alter the model by allowing a funding period, such that abatement costs become effective after a period $\Delta T_{\mathrm{A}}$. For the funding, interest has to be paid such that
$C_{\mathrm{A}}(t+\Delta T_{A})\ :=\ C_{\mu}(t)\ \left(1+FR(t,t+\Delta T_{A};t)\Delta T_{\mathrm{A}}\right).$ (4)
@@
For the case of no funding, i.e. $\Delta T_{\mathrm{A}}=0$, $C_{\mu}=C_{\mathrm{A}}$, and we recover the original model.
@@
Since damages occur instantaneously, we do not consider funding of these. However, we consider a kind of non-linear discounting.
Non-Linear Discounting of Damage Cost
The idea behind non-linear discounting is that the discount factor depends on the level of the amount to be discounted. This can be justified by a model where one has to attribute for the higher default risk of excessively large future cash-flows. The mechanics are that the compensation of this default risk lowers the return (the interest rate) and hence increases the discount factor38.
Put differently, this adds an increasing weight to large or very large damages.
@@
Let $C^{\circ}_{\mathrm{D}}(t)$ denote the damage cost of the classical model, i.e. the quantity that was formerly denoted by $C_{\mathrm{D}}(t)$. We then re-define the effective damage costs as
@@
where $DC(C^{\circ}_{\mathrm{D}}(t);t)$ is the default compensation factor. It is somewhat similar to the inverse of a discount factor, describing the over-proportional cost of funding large projects.
In our experiment, we can specify this weight as a function of the damage $C_{\circ}(t)$ or as a function of the fraction of the GDP, i.e., $C_{\circ}(t)/GDP(t)$.
The latter approach allows us to penalise damages that exceed a certain percentage of the GDP.
@@
The model code for the stochastic DICE model including the proposed extensions, as well as the experiments generating the data and figures is available in the following Git repository https://gitlab.com/finmath/finmath-climate-nonlinear.
@@ Acknowledgements
C.F. and L.Q. developed the research idea together. C.F. lead the re-implementation of the DICE model and extensions as well as the experiments with input from L.Q. L.Q. lead the plotting implementation with input from C.F. C.F. and L.Q. analysed and interpreted results. C.F. derived the analytic expression for generational inequality. C.F and L.Q. wrote the manuscript.
@@
Accounting for Financing Risks improves
Intergenerational Equity of Climate Change Mitigation
Christian P. Fries and Lennart Quante
@@
Supplementary Figure 1: The sensitivity $\frac{\mathrm{d}C_{\mathrm{A}}(t)}{\mathrm{d}T^{\mu=1}}$ and $\frac{\mathrm{d}C_{\mathrm{D}}(t)}{\mathrm{d}T^{\mu=1}}$ of abatement cost $C_{\mathrm{A}}$ and damage cost $C_{\mathrm{D}}$ on changes of the abatement policy $T^{\mu=1}$, without weighting the values by $\frac{\mathrm{d}V(t)}{\mathrm{d}C(t)}\frac{N(0)}{N(t)}$. This is the nominal sensitivity to abatement policy changes, entering the calibration.
Supplementary Figure 2: Deterministic model: Financing extensions lead to faster abatement compared to standard model a emissions, b abatement, c damages, and d total GDP relative cost
supfigure]supfig:pathways_deterministic_with_extension
Supplementary Figure 3: Deterministic model: Total cost relative to GDP by year of birth in the different model specifications up to 2100 Normalised by annual GDP and scaled such that blue means lower, red higher average total cost values than 3% of GDP for this generation.
supfigure]supfig:generational_burden_deterministic
Supplementary Figure 3: Deterministic model: Total cost relative to GDP by year of birth in the different model specifications up to 2100 Normalised by annual GDP and scaled such that blue means lower, red higher average total cost values than 3% of GDP for this generation.
Supplementary Figure 2: Deterministic model: Financing extensions lead to faster abatement compared to standard model a emissions, b abatement, c damages, and d total GDP relative cost
Supplementary Figure 1: The sensitivity $\frac{\mathrm{d}C_{\mathrm{A}}(t)}{\mathrm{d}T^{\mu=1}}$ and $\frac{\mathrm{d}C_{\mathrm{D}}(t)}{\mathrm{d}T^{\mu=1}}$ of abatement cost $C_{\mathrm{A}}$ and damage cost $C_{\mathrm{D}}$ on changes of the abatement policy $T^{\mu=1}$, without weighting the values by $\frac{\mathrm{d}V(t)}{\mathrm{d}C(t)}\frac{N(0)}{N(t)}$. This is the nominal sensitivity to abatement policy changes, entering the calibration.

### v2

@@ 4.2.2 Sensitivity of Damage and Abatement Cost to Policy Changes
From eq. 6 we see why the model calibration may result in some intergenerational inequality: Let $s$ be a fixed time at which a change in the abatement policy $\mu(s)$ is considered. Since $\mathrm{d}C_{\mathrm{D}}/\mathrm{d}\mu(s)>0$ and $\mathrm{d}C_{\mathrm{A}}/\mathrm{d}\mu(s)<0$ we see from eq. 6 that the model equates the increment of the cost of damage to the increment of the cost of abatement, both weighted by the value-to-cost-sensitivity $\mathrm{d}V(t)/\mathrm{d}C(t)\cdot N(0)/N(t)$. We have
$\int_{0}^{T}\frac{\mathrm{d}V(t)}{\mathrm{d}C(t)}\frac{\mathrm{d}C_{\mathrm{D}}(t)}{\mathrm{d}\mu(s)}\ \frac{N(0)}{N(t)}\mathrm{d}t\ +\ \int_{0}^{T}\frac{\mathrm{d}V(t)}{\mathrm{d}C(t)}\frac{\mathrm{d}C_{\mathrm{A}}(t)}{\mathrm{d}\mu(s)}\ \frac{N(0)}{N(t)}\mathrm{d}t\ \stackrel{{\scriptstyle!}}{{=}}\ 0\text{,}$ (7)
where the to cost function $t\mapsto\frac{\mathrm{d}C_{\mathrm{A}}(t)}{\mathrm{d}\mu(s)}$ and $t\mapsto\frac{\mathrm{d}C_{\mathrm{D}}(t)}{\mathrm{d}\mu(s)}$ have different temporal distributions.
A change in the abatement policy in time $s$ affects the abatement cost $C_{\mathrm{A}}(s)$ only at time $s$, whereas it affects damage cost $C_{\mathrm{D}}(t)$ for possibly all $t\geq s$. This motivates to consider the damage cost per abatement cost induced by a change in the abatement policy $\mu(s)$, i.e., $\frac{\partial C_{\mathrm{D}}(t)}{\partial\mu(s)}\big{/}\frac{\partial C_{\mathrm{A}}(s)}{\partial\mu(s)}$. As the model calibration is weighting cost by the cost-to-value weight, we consider the corresponding weighted damage-cost per abatement-cost sensitivity. The minus in (8) just accounts for the fact that abatement cost sensitivity and damage cost sensitivity have opposite signs.
$\rho_{\mathrm{D}}(s;t)\ :=\ -\frac{\frac{\partial V(t)}{\partial C(t)}\cdot\frac{\partial C_{\mathrm{D}}(t)}{\partial\mu(s)}\cdot\frac{N(0)}{N(t)}}{\frac{\partial V(s)}{\partial C(s)}\cdot\frac{\partial C_{\mathrm{A}}(s)}{\partial\mu(s)}\cdot\frac{N(0)}{N(s)}}\text{.}$ (8)
From Equation 5 we see that for fixed $s$ the integral over $t\mapsto\rho(s;t)$ equals $1$, i.e., $\rho_{\mathrm{D}}(s)$ can be interpreted as a density. Hence, we may interpret $\int_{0}^{\infty}t\cdot\rho_{\mathrm{D}}(s;t)\ \mathrm{d}t$ as the expected time of damage cost per avoided abatement cost in time $s$.
Figure 8: The density $t\mapsto\rho_{\mathrm{D}}(s;t)$ for $s=10,20,50$.
In the classical DICE 2016 model the expected time of damage cost per avoided abatement cost is approximately 70 years after the time of abatement (Figure 8), indicating that balancing marginal abatement and damage cost is a matter of intergenerational equity.
Sensitivity of Cost to Time of Reaching 100 % Abatement
The sensitivity $\frac{\mathrm{d}}{\mathrm{d}\mu}$ is a functional derivative or gradient. To visualize the aggregated dependency structure, we consider the one-parameter deterministic abatement model (4). For this abatement model, the equilibrium state is given by
[…]
Given the one-parameter model, we may depict the (weighted) damage cost sensitivity $\frac{\mathrm{d}C_{\mathrm{D}}(t)}{\mathrm{d}T^{\mu=1}}$ and the (weighted) abatement cost sensitivity $\frac{\mathrm{d}C_{\mathrm{A}}(t)}{\mathrm{d}T^{\mu=1}}$, see fig. 2.
4.2.3 Cost per GDP
To measure the burden born by a generation, we consider the relative total damage and abatement cost per GDP. Putting nominal values relative to the GDP is a common measure, e.g., to assess national debt [45]. Here, it reflects that a more wealthy generation can carry a greater burden, mimicking an effort sharing scheme and also considering the greater capacity of future generations under the expected baseline growth in DICE.
In our numerical experiments we measure the instantaneous cost-per-GDP $C(s)/GDP(s)$ as well as the lifetime-average cost-per-GDP $\overline{C}(s)/\overline{GDP}(s)$.
To allow for a netting inside a generation one may consider the (discounted) average value over a generation lifetime. For a given time $s$ let $T^{\mathrm{L}}(s)$ denote the expected lifetime of a generation. We then consider the lifetime average cost and GDP.
$\overline{C}(s)\ :=\ \int_{s}^{s+T^{\mathrm{L}}(s)}C(t)\frac{N(s)}{N(t)}\mathrm{d}t\text{,}\qquad\overline{GDP}(s)\ :=\ \int_{s}^{s+T^{\mathrm{L}}(s)}GDP(t)\frac{N(s)}{N(t)}\mathrm{d}t\text{.}$
4.3 Model Extensions
We present modifications to integrated assessment models to investigate the effect of stochastic interest rates, funding of abatement costs, and non-linear financing cost [41]. These may be also incorporated into more complex IAMs. We use the simple DICE model to illustrate potential effects.
4.3.1 Stochastic Interest Rates
The classical way in which interest rates are modelled in an IAM is via a discount factor. A time-$t$ value $V(t)$ is derived in the model, then discounted and aggregated to a final value forming the objective function
$\int_{0}^{T}V(t)\exp\left(-\int_{0}^{t}r(s)\mathrm{d}s\right)\mathrm{d}t\text{.}$
We model stochastic interest rates $r$. Our implementation allows to use a general discrete forward rate model (LIBOR Market Model). Our results were conducted with a classical Hull-White model. The model provides the stochastic numeráire $N$ (and thus discount factor) as well as the stochastic forward rate $FR$. Setting the volatility of the interest rate to zero recovers the deterministic model with a possibly time-dependent interest rate (term-structure of interest rates). For exemplary parameterization, we set the mean reversion speed to $a(t)=0.02$ and the volatility to $\sigma(t)=0.4\%$. While our result should be interpreted as a qualitative result, showing that stochastic interest rates generate adverse effects, the interest rate scenarios generated by our parameter choice are realistic, rather conservative: the 1st and 99th percentile of the generated interest rates at t = 100 years are -1 % and 5 %, respectively. For a corresponding visualization, see [39].
If interest rates are stochastic but the function $V(t)$ remains deterministic, adding stochastic interest rates does not change the interaction with the IAM since
$\begin{split}\mathrm{E}\left(\int_{0}^{T}V(t)\exp\left(-\int_{0}^{t}r(s)\mathrm{d}s\right)\mathrm{d}t\right)&\ =\ \int_{0}^{T}V(t)\mathrm{E}\left(\exp\left(-\int_{0}^{t}r(s)\mathrm{d}s\right)\right)\mathrm{d}t\\ &\ =\ \int_{0}^{T}V(t)\exp\left(-\int_{0}^{t}\bar{r}(s)\mathrm{d}s\right)\mathrm{d}t\text{,}\end{split}$ (10)
with $\bar{r}(t)\ =\ -\frac{\partial}{\partial t}\log\left(\mathrm{E}\left(\exp\left(-\int_{0}^{t}r(s)\mathrm{d}s\right)\right)\right)$.
Due to the lack of a feedback that allows to adjust the abatement path and resulting damages, adding stochastic interest rates does not yet introduce changes in the model dynamics. This changes however, if one assumes that the abatement policy is adapted to the stochastic changes, here the change in interest rate level, i.e., if one allows $\mu$ to be an (adapted) stochastic process. [^2] Our choice to make interest rates stochastic is only exemplary. The interest rate is an important factor in linking present abatement cost to the avoided future damage cost. While one may introduce stochasticity in many other state variables, the use of an abatement policy adapted to the economic factors (interest rates) will already lead to all other state variables becoming stochastic.
We use a classical Hull-White model, that is
$\begin{split}\mathrm{d}r(t)&\ =\ (\theta(t)-a(t)r(t))\ \mathrm{d}t\ +\ \sigma(t)\ \mathrm{d}W(t)\text{,}\quad r(t_{0})\ :=\ r_{0}\text{,}\\ \mathrm{d}N(t)&\ =\ r(t)\ \mathrm{d}t\text{,}\end{split}$ (11)
where an exact time-discretization scheme [47] was used. The model’s mean-reversion level $\theta$ is such that the numéraire-relative zero coupon bond prices are martingales and their expectation agrees with the classical non-stochastic model. In other words, our stochastic interest rate model is such that the optimal deterministic abatement strategy would result in the same optimal emission pathway as the model with deterministic interest rates.
4.3.2 Stochastic Abatement Policy
Assuming that a change in the interest rates triggers a re-calibration of the model, a stochastic interest rate implies a stochastic abatement policy, i.e., $\mu$ will become stochastic. The abatement policy $t\mapsto\mu(t)$ becomes a stochastic process that adapts to the changes in interest rates. This reflects the possibility that the abatement policy can be adjusted to the interest rate scenarios.
The determination of an optimal stochastic abatement policy is a classical optimal exercise problem, as in the valuation of financial products with early exercise options in mathematical finance. Here, $\mu$ takes the role of the exercise strategy. As the time-$t$ optimal exercise is required to be a $\mathcal{F}_{t}$-measurable, it may be represented as a functional form of time-$t$ measurable random variables, see [46].
An approach to determine the optimal exercise strategy is to parameterize $\mu(t)$ as a functional form of $\mathcal{F}_{t}$-measurable random variables and apply a global optimization. As we utilize a one-factor Markovian model for the short rate $r$, we consider a functional form where $\mu$ depends on $t$ and the current interest rate level $r(t)$.
A simple example of a stochastic abatement model is a parametric one, where the abatement speed is a (linear) function of the interest rate level $r$, e.g.,
@@ 4.3.2 Stochastic Abatement Policy
The simplicity of the models allows us to interpret the parameters. We have that $a_{0}$ is the abatement speed at the zero interest rate level and $a_{1}$ describes the change of the abatement speed by interest rate change,
$a_{0}\ =\ \frac{\partial\mu(t)}{\partial t}\big{|}_{r=0}\text{,}\qquad a_{1}\ =\ \frac{\partial}{\partial r}\ \frac{\partial\mu(t)}{\partial t}\text{.}$
@@
Obviously, the stochastic abatement models may achieve a larger-or-equal welfare or lower-or-equal welfare-risk than the corresponding deterministic model (4) simply due to the additional degree of freedom $a_{1}$ in the abatement optimisation. The calibration of the optimal model (12) results in a negative value for $a_{1}$, reflecting that higher interest rate correspond to slower abatement speed.
4.3.3 Funding of Abatement Costs
The geophysical part of the model models a global mean temperature level, which results from past emissions. The simple damage function of the DICE model provides damages as a reduction of global GDP. Damages may be reduced by performing abatement of emissions. Both abatement and damage are associated with costs. The time $t_{i}$ costs are deduced from the time $t_{i}$ GDP. The remainder is then available for consumption or investment. Investment adds to the capital which determines the GDP of the next time $t_{i+1}$. However, abatement cost and damage cost linked to the same abatement policy decision occur at significantly different times.
We introduce the option of funding of abatement cost. As abatement of emissions is a planned process of societal relevance, it is reasonable to assume that abatement cost are covered by a loan for which interest rate corresponds to the current discount rate. Our model allows to apply a funding spread, i. e. higher interest rates for loans than for discounting, but this is not considered in this general introduction of the phenomenon. In future work, the concept of the climate beta could be used to introduce an interest rate spread for abatement investments, as introduced in DICE-2023 [38].
We define $C_{\mu}(t)$ as the instantaneous abatement costs in time $t$, i.e., the quantity that was formerly denoted by $C_{\mathrm{A}}(t)$. We allow that $C_{\mu}(t)$ might be funded for a period $\Delta T_{\mathrm{A}}$. Thus abatement costs of $C_{\mu}(t)$ are accrued with the forward rate $FR(t,t+\Delta T_{\mathrm{A}};t)$ observed in $t$, such that the realized abatement cost $C_{\mathrm{A}}$ at time $t+\Delta T_{\mathrm{A}}$ are
$C_{\mathrm{A}}(t+\Delta T_{A})\ :=\ C_{\mu}(t)\ \left(1+FR(t,t+\Delta T_{A};t)\Delta T_{\mathrm{A}}\right)\text{.}$ (13)
@@ 4.3.3 Funding of Abatement Costs
For the case of no funding, i.e. $\Delta T_{\mathrm{A}}=0$, we recover the classical model with $C_{\mu}=C_{\mathrm{A}}$.
In a standard model for risk-neutral valuation, this change would have no effect, as the accruing is compensated by the discounting. However, the change might introduce an effect in the DICE model, due to the way how abatement and damage cost are associated and due to the time-preference included in the utility function.
@@
Since damages occur instantaneously, we do not consider funding of these. The total cost is given by $C(t):=C_{\mathrm{A}}(t)+C_{\mathrm{D}}(t)$.
4.3.4 Non-Linear Financing Costs
For an unsecured financial cash-flow its present value is defined by a discount factor times the cash-flow. If the future cash-flow is subject to default, the discount factor is lower, reflecting the additional value reduction due to the risk of (partial) default. As default is not an option for future damage cost, it appears as if the risk-free discount factor should apply. However, since no hedging strategy exists, a risk free funding is not possible. Thus additional cost may occur to secure the unsecured funding [41]. Since damage cost may become very large - much larger than funds provided by standard financial markets - it is reasonable to assume that these additional funding cost become (over-proportionally) large for larger cost. We model this by optionally adding non-linear financing cost using a non-linear funding model [41].
This model allows, that the discount factor may depend on the magnitude of the cash-flow. Thus the funding of larger cash-flows requires a premium to compensate for a larger default risk or other frictions. For our application, this means that larger cost get a larger weight.
@@ 4.3.4 Non-Linear Financing Costs
Our model modification is now a modification of the damage cost. Let $C^{\circ}_{\mathrm{D}}(t)$ denote the damage cost of the classical model, i.e. the quantity that was formerly denoted by $C_{\mathrm{D}}(t)$. We then re-define the effective damage costs as
@@
where $DC(C^{\circ}_{\mathrm{D}}(t);t)$ is the default compensation factor. It is somewhat similar to the inverse of a discount factor, describing the over-proportional cost to fund large projects.
As $C^{\circ}_{\mathrm{D}}(t)$ represents a time-value, it is natural that the default compensation factor depends on a normalized value only. We allow two different normalizations: either with the numéraire $N(t)$ or with the GDP $Y(t)$. Hence, our model for the default compensator is
$DC(C^{\circ}_{\mathrm{D}}(t);t)\ =\ DC^{N}(C^{\circ}_{\mathrm{D}}(t);t)\ :=\ DC^{*}(\frac{C^{\circ}_{\mathrm{D}}(t)}{N(t)})\text{,}$ (14)
or, alternatively,
$DC(C^{\circ}_{\mathrm{D}}(t);t)\ =\ DC^{Y}(C^{\circ}_{\mathrm{D}}(t);t)\ :=\ DC^{*}(\frac{C^{\circ}_{\mathrm{D}}(t)}{Y(t)})\text{.}$ (15)
The latter approach allows us to penalize damages that exceed a certain percentage of the GDP.
The factor $DC^{*}(x)$ depends on the size of $x$. For small $x$ we have $DC^{*}(x)=1$, but for large $x$ we may have factors $>1$. Obviously, this will penalize large spikes in the costs.
In our numerical experiments the funding period was set to 20 years and the non-linear discounting was set to increase the default compensator significantly for damage cost over 3 % of the GDP (black). The parameter choice is exemplary, other experiments with similar results can be reproduced in the published source code.
As the abatement cost are usually smaller and part of a more planned process, we do not consider a default compensation factor for the abatement cost.
@@
The model code for the stochastic DICE model, including the proposed extensions, as well as the experiments generating the data and figures, is available in the following Git repository https://gitlab.com/finmath/finmath-climate-nonlinear.
@@ Acknowledgements
C.F. and L.Q. developed the research idea together. C.F. lead the re-implementation of the DICE model and extensions, as well as the experiments with input from L.Q. L.Q. lead the plotting implementation with input from C.F. C.F. and L.Q. analyzed and interpreted results. C.F. derived the analytic expression for generational inequality. C.F and L.Q. wrote the manuscript.
@@
Intergenerational Equitable Climate Change Mitigation:
Negative Effects of Stochastic Interest Rates;
Positive Effects of Financing
[…]
@@
Supplementary Figure 1: Increasing interest rate level increases the time of 100% abatement $T^{\mu=1}$.
Supplementary Figure 2: Increasing interest rate volatility, i.e. riskiness, decreases the time of 100% abatement $T^{\mu=1}$.